# HalluGuard: Demystifying Data-Driven and Reasoning-Driven Hallucinations in LLMs

### v1

@@ B.1 Preliminaries and Notation
Let $\mathcal{K}\in\mathbb{R}^{r\times r}$ be the NTK Gram matrix formed on $r$ light semantic perturbations (see Assumptions A1–A4 in the main theory section). Denote its eigen decomposition by $\mathcal{K}=V\Lambda V^{\top}$ with
@@ B.2 Representational Adequacy via $\det(\mathcal{K})$ with Explicit Constants
    (Source condition) There exist $s>0$ and $R_{s}>0$ such that $u^{*}\in\mathrm{Range}(\Lambda^{s})$, i.e., $\sum_{i=1}^{r}\frac{\langle u^{*},v_{i}\rangle^{2}}{\lambda_{i}^{2s}}\leq R_{s}^{2}$. This is standard in kernel approximation and encodes RKHS regularity.
@@
    (Spectral envelope) There exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ and $\alpha>1$ such that $\lambda_{i}\leq\overline{\lambda}$ for all $i$ and $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$. (Polynomial decay is a common stylization; other envelopes can be treated similarly.)
@@
Write $u^{*}=\sum_{i\geq 1}c_{i}v_{i}$ with $c_{i}=\langle u^{*},v_{i}\rangle$. Then $\|u^{*}-\Pi_{U_{h}}u^{*}\|^{2}=\sum_{i>r}c_{i}^{2}\leq\sum_{i>r}\lambda_{i}^{2s}\cdot\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}\sum_{i>r}\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}R_{s}^{2}.$
To connect $\lambda_{r+1}$ (or $\lambda_{r}$) to $\det(\mathcal{K})$, we need an explicit lower bound of the form $\lambda_{r}\geq\underline{c}\,\det(\mathcal{K})^{\,\theta}$ with constants $(\underline{c},\theta)$ depending on the spectral envelope. The following inequality suffices.
@@
Under A1–A3 and S1S2,
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\;\det(\mathcal{K})^{-\,c_{d}}\;\|u^{*}\|,$
with
$c_{d}\;=\;\frac{s}{\,r-1\,}\quad\text{and}\quad C_{d}\;=\;\overline{\lambda}^{\,s}\,\frac{R_{s}}{\|u^{*}\|}.$
@@
By Lemma B.1 with $\lambda_{r+1}\leq\lambda_{r}$, $\inf_{u\in U_{h}}\|u^{*}-u\|\leq R_{s}\,\lambda_{r}^{\,s}.$ Lemma B.2 gives $\lambda_{r}^{\,s}\geq\det(\mathcal{K})^{\,s}/\overline{\lambda}^{\,s(r-1)}$; rearranging,
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;R_{s}\,\overline{\lambda}^{\,s(r-1)}\,\det(\mathcal{K})^{-\,s}.$
Rescale constants relative to $\|u^{*}\|$ by setting $C_{d}:=\overline{\lambda}^{\,s}\,(R_{s}/\|u^{*}\|)$ and $c_{d}:=s/(r-1)$ to obtain the stated form:
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;\big(\overline{\lambda}^{\,s}\,\tfrac{R_{s}}{\|u^{*}\|}\big)\;\det(\mathcal{K})^{-\,s/(r-1)}\;\|u^{*}\|.$
The variant using the envelope $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$ is obtained by combining $\det(\mathcal{K})\leq\overline{\lambda}^{\,r-1}\lambda_{r}$ with the explicit lower bound on $\lambda_{r}$, yielding the alternative exponent shown. ∎
Numerical note (stable surrogate).
In practice we use $\log\det(\mathcal{K})$ via Cholesky and aggregate with $z$-normalization across components to avoid scale domination by any single term.
@@ C.2 Ablation Study on $\log\kappa^{2}$
C.3 Computational Efficiency Analysis
To assess practical deployment feasibility, we measured inference latency on an NVIDIA A100/H200 GPU. Our setup utilizes batched parallel sampling to generate $K=10$ trajectories, ensuring sub-linear scaling of the computational cost. The core HalluGuard operations—specifically feature clipping and computing the NTK score via the Gram matrix—add minimal latency, requiring less than 1 ms of post-processing time per query.
@@ C.3 Computational Efficiency Analysis
C.4 Detection Performance Analysis
@@
RAGTruth SAPLMA 72.80 70.10 72.20 63.50 55.10 78.90 74.20 74.10 68.00 58.20 79.40 77.30 79.00 72.10 60.50 81.00 78.20 79.44 72.80 61.30 74.20 72.10 70.50 61.80 55.90
RAGTruth LLM-Check 68.10 64.50 63.90 55.20 44.80 72.30 68.40 66.50 57.90 46.30 75.20 71.60 67.40 60.30 48.70 76.10 73.20 68.90 61.10 49.50 71.60 68.90 63.20 55.40 46.10
RAGTruth ITI 69.30 65.80 66.10 57.90 47.90 73.10 69.20 68.20 59.80 49.10 76.00 72.50 69.40 61.80 50.90 77.20 74.10 70.50 62.40 51.70 72.80 70.10 65.40 57.10 47.80
@@
HaluEval SAPLMA 67.10 63.20 69.20 62.10 54.00 69.50 65.70 68.30 61.60 53.20 72.00 68.40 75.10 69.30 58.90 71.20 68.10 75.40 70.30 58.50 76.10 72.20 76.80 70.60 60.90
HaluEval LLM-Check 63.50 59.40 61.10 53.00 44.50 66.80 62.90 65.40 57.50 47.50 70.10 66.30 63.80 57.20 47.10 69.30 65.40 66.20 59.50 49.00 71.50 67.60 63.50 55.90 47.40
HaluEval ITI 64.80 60.70 63.40 55.20 46.80 67.40 63.50 66.90 58.60 49.40 71.00 67.20 66.10 59.10 48.60 70.20 66.30 68.10 61.10 50.60 72.30 68.20 65.20 57.50 48.70
C.5 Tightness of Bound
@@
C.6 Correlation of reasoning-driven and data-driven terms with different types of datasets
@@
C.7 Case Study
Case Study 1 — GSM8K (Multi-step Arithmetic): Bias $\rightarrow$ Drift $\rightarrow$ Snowballing.
@@ C.7 Case Study
Case Study 2 — Long-Document Summarization: Misalignment $\rightarrow$ Overreach $\rightarrow$ Fabrication.
@@ Appendix D Usage of LLM
Large language models (LLMs) were employed in a limited and transparent manner during the preparation of this manuscript. Specifically, LLMs were used to assist with linguistic refinement, style adjustments, and minor text editing to improve clarity and readability. They were not involved in formulating the research questions, designing the theoretical framework, conducting experiments, or interpreting results. All scientific contributions—including conceptual development, methodology, analyses, and conclusions—are the sole responsibility of the authors.

### v2

@@ B.1 Preliminaries and Notation
Let $\mathcal{K}\in\mathbb{R}^{r\times r}$ be the NTK Gram matrix formed on $r$ light semantic perturbations (see Assumptions A1-A3 in the main theory section). Denote its eigen decomposition by $\mathcal{K}=V\Lambda V^{\top}$ with
@@ B.2 Representational Adequacy via $\det(\mathcal{K})$ with Explicit Constants
    (Source condition) Let $\mathcal{T}$ denote the infinite-dimensional NTK integral operator. We assume there exists a regularity exponent $s>0$ and a constant $R_{s}>0$ such that $u^{*}\in\mathrm{Range}(\mathcal{T}^{s})$. Equivalently, the spectral coefficients satisfy: $\sum_{i=1}^{r}\frac{\langle u^{*},v_{i}\rangle^{2}}{\lambda_{i}^{2s}}\leq R_{s}^{2}$. This is standard in kernel approximation and encodes RKHS regularity.
@@
    (Spectral envelope) Let $\overline{\lambda}$ and $\underline{\lambda}$ denote uniform upper and lower bounds on the kernel spectrum. We assume there exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ and a decay rate $\alpha>1$ such that $\lambda_{i}\leq\overline{\lambda}$ for all $i$, and the tail eigenvalue satisfies $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$. (Polynomial decay is a common stylization; other envelopes can be treated similarly.)
@@
Write $u^{*}=\sum_{i\geq 1}c_{i}v_{i}$ with $c_{i}=\langle u^{*},v_{i}\rangle$. By the source condition, $\|u^{*}-\Pi_{U_{h}}u^{*}\|^{2}=\sum_{i>r}c_{i}^{2}\leq\sum_{i>r}\lambda_{i}^{2s}\cdot\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}\sum_{i>r}\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}R_{s}^{2}.$
To connect the representation error to the empirical NTK Gram matrix $\mathcal{K}$, we leverage the algebraic relationship between the smallest eigenvalue $\lambda_{r}$ and the determinant.
@@
Under A1-A3 and S1-S2, Under A1–A3 and S1S2, the approximation error is bounded by:
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\;\det(\mathcal{K})^{\,c_{d}},$
with explicit constants independent of the target sequence:
$c_{d}\;=\;\frac{s}{r}\quad\text{and}\quad C_{d}\;=\;R_{s}.$
@@
By Lemma B.1 and the fact that eigenvalues are monotonically decreasing ($\lambda_{r+1}\leq\lambda_{r}$), we have:
$\inf_{u\in U_{h}}\|u^{*}-u\|\leq R_{s}\,\lambda_{r}^{\,s}.$
Recall that the determinant of the empirical Gram matrix is the product of its eigenvalues, $\det(\mathcal{K})=\prod_{i=1}^{r}\lambda_{i}$. Since $\lambda_{r}$ is the minimum eigenvalue of the rank-$r$ matrix, it follows strictly that $\lambda_{r}^{r}\leq\det(\mathcal{K})$, which implies $\lambda_{r}\leq\det(\mathcal{K})^{1/r}$. Raising both sides to the power of $s$ yields $\lambda_{r}^{s}\leq\det(\mathcal{K})^{s/r}$. Substituting this upper bound into the approximation error gives:
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;R_{s}\det(\mathcal{K})^{\,s/r}.$
Setting $C_{d}:=R_{s}$ and $c_{d}:=s/r$ completes the proof. ∎
In practice, tracking the direct determinant can cause numerical underflow in high-dimensional spaces. We use $\log\det(\mathcal{K})$ via the Cholesky decomposition as our empirical score, aggregating with $z$-normalization across components to avoid scale domination by any single dimension.
@@ C.2 Ablation Study on $\log\kappa^{2}$
Table 6: Ablation on stability term $\log\kappa^{2}$ (MATH500).
Method Pearson $R$ MSE
HalluGuard 0.985 0.0192
w/o $\log\kappa^{2}$ 0.8904 0.0381
The error in table 6 nearly doubles without the stability term, confirming that spectral conditioning is essential for stable reasoning-risk quantification.
C.3 Ablation Study on Semantic Encoder $\Phi$
To examine sensitivity to the semantic encoder $\Phi$, we replace the default representation with widely adopted alternatives, including BERT, SimCSE, and E5. We evaluate across multiple backbone models and benchmarks.
Table 7 reports AUROC and AUPRC on RAGTruth, GSM8K, and TruthfulQA. Across all settings, HalluGuard consistently outperforms encoder-substituted variants. For example, on QwQ-32B (RAGTruth), replacing the default encoder with BERT reduces AUROC from 84.59 to 81.44.
These results indicate that the performance gain does not stem from surface semantic similarity of final outputs. Instead, the method captures geometric structure of reasoning trajectories, which external encoders cannot fully preserve.
Table 7: Encoder ablation across backbones and benchmarks (AUROC / AUPRC).
Backbone Method RAGTruth GSM8K TruthfulQA AUROC AUPRC AUROC AUPRC AUROC AUPRC GPT-2 HalluGuard 75.51 73.40 72.04 69.88 72.10 68.76 +BERT 72.48 70.12 67.31 64.90 68.02 65.01 +SimCSE 73.21 71.05 68.44 66.02 69.14 66.27 +E5 74.02 71.66 69.12 66.80 70.03 67.10 OPT-6.7B HalluGuard 80.13 76.77 72.57 70.31 69.59 68.36 +BERT 77.44 74.20 67.95 65.48 66.12 64.80 +SimCSE 78.11 74.83 69.01 66.40 67.08 65.72 +E5 78.66 75.31 70.04 67.25 67.80 66.41 Mistral-7B HalluGuard 82.31 80.79 80.62 77.30 77.05 73.79 +BERT 79.02 76.91 75.51 72.08 73.14 69.52 +SimCSE 79.88 77.66 76.40 73.01 74.08 70.40 +E5 80.41 78.20 77.12 73.74 74.66 71.05 QwQ-32B HalluGuard 84.59 81.15 75.81 74.68 74.26 72.76 +BERT 81.44 78.03 70.92 68.90 70.35 68.01 +SimCSE 82.10 78.66 72.10 69.82 71.20 68.70 +E5 82.66 79.12 73.05 70.44 72.02 69.31 LLaMA2-13B HalluGuard 77.51 75.30 79.01 76.73 78.50 77.56 +BERT 74.26 72.04 73.12 70.60 74.41 72.88 +SimCSE 75.11 72.83 74.20 71.51 75.36 73.54 +E5 75.78 73.44 75.14 72.32 76.10 74.22
C.4 Computational Efficiency Analysis
To assess practical deployment feasibility, we measured inference latency on an NVIDIA A100/H200 GPU. Our setup utilizes batched parallel sampling to generate $K=10$ trajectories, ensuring sub-linear scaling of the computational cost. The core HalluGuard operations-specifically feature clipping and computing the NTK score via the Gram matrix-add minimal latency, requiring less than 1 ms of post-processing time per query.
@@ C.4 Computational Efficiency Analysis
C.5 Detection Performance Analysis
@@
SAPLMA 72.80 70.10 72.20 63.50 55.10 78.90 74.20 74.10 68.00 58.20 79.40 77.30 79.00 72.10 60.50 81.00 78.20 79.44 72.80 61.30 74.20 72.10 70.50 61.80 55.90
LLM-Check 68.10 64.50 63.90 55.20 44.80 72.30 68.40 66.50 57.90 46.30 75.20 71.60 67.40 60.30 48.70 76.10 73.20 68.90 61.10 49.50 71.60 68.90 63.20 55.40 46.10
ITI 69.30 65.80 66.10 57.90 47.90 73.10 69.20 68.20 59.80 49.10 76.00 72.50 69.40 61.80 50.90 77.20 74.10 70.50 62.40 51.70 72.80 70.10 65.40 57.10 47.80
@@
SAPLMA 67.10 63.20 69.20 62.10 54.00 69.50 65.70 68.30 61.60 53.20 72.00 68.40 75.10 69.30 58.90 71.20 68.10 75.40 70.30 58.50 76.10 72.20 76.80 70.60 60.90
LLM-Check 63.50 59.40 61.10 53.00 44.50 66.80 62.90 65.40 57.50 47.50 70.10 66.30 63.80 57.20 47.10 69.30 65.40 66.20 59.50 49.00 71.50 67.60 63.50 55.90 47.40
ITI 64.80 60.70 63.40 55.20 46.80 67.40 63.50 66.90 58.60 49.40 71.00 67.20 66.10 59.10 48.60 70.20 66.30 68.10 61.10 50.60 72.30 68.20 65.20 57.50 48.70
C.6 Tightness of Bound
@@
C.7 Correlation of reasoning-driven and data-driven terms with different types of datasets
@@
C.8 Case Study
Case Study 1 - GSM8K (Multi-step Arithmetic): Bias $\rightarrow$ Drift $\rightarrow$ Snowballing.
@@ C.8 Case Study
Case Study 2 - Long-Document Summarization: Misalignment $\rightarrow$ Overreach $\rightarrow$ Fabrication.
@@
C.9 Comparison with Inside and MIND
Inside and MIND serve as empirical uncertainty diagnostics. Inside analyzes covariance spectra of static representations, while MIND measures temporal variations in hidden states. Both methods extract post-hoc signals and produce a single uncertainty score.
In contrast, HalluGuard derives a structured risk decomposition from generative dynamics, separating data-driven and reasoning-driven sources via NTK spectral geometry and instability amplification. This formulation explicitly models compounded reasoning errors.
We evaluate all methods on the Snowballing benchmark Zhang et al. (2023), which emphasizes progressive reasoning instability. As shown in Table 12, HalluGuard consistently outperforms Inside and MIND across all backbone models.
Table 12: Comparison with Inside and MIND on the Snowballing benchmark across different backbone models (AUROC / AUPRC).
Method GPT-2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-7B LLaMA2-70B HalluGuard 88.52/82.14 92.63/87.42 94.87/89.66 97.41/95.08 93.28/88.03 97.96/95.37 Inside 74.11/66.39 78.24/70.51 83.32/75.80 87.55/80.47 81.72/73.11 89.03/82.77 MIND 69.42/58.73 74.56/64.37 78.67/68.52 84.03/73.68 77.91/65.89 86.28/78.41
C.10 Additional Evaluation on Multimodal and Long-Context Regimes
To evaluate generalization beyond short-form reasoning tasks, we extend experiments to (i) multimodal hallucination detection on POPE Li et al. (2023b), and (ii) long-context generation on GovReport Huang et al. (2021) and NarrativeQA s Kočiský et al. (2018).
Across all backbone models, HalluGuard consistently achieves the strongest AUROC and AUPRC, demonstrating robustness under multimodal noise and long-range dependency drift.
Multimodal Hallucination (POPE).
Table 13: Comparison of methods across different backbone models on POPE(AUROC/AUPRC).
Method GPT-2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-7B LLaMA2-70B Perplexity 61.12/53.04 68.27/60.18 72.41/64.09 79.36/73.22 70.15/62.31 83.48/76.19 HalluGuard 74.33/68.27 81.22/75.36 86.47/80.51 91.58/86.42 85.39/78.44 94.63/89.27 Inside 70.08/64.12 77.19/70.33 83.44/75.28 89.27/82.36 81.22/74.41 92.51/87.39 MIND 66.17/58.22 73.31/66.14 79.28/71.39 86.44/79.33 77.18/69.27 89.36/83.48 LN-Entropy 63.09/55.11 71.24/62.18 76.37/67.06 84.33/75.29 74.12/65.18 87.42/80.33 Energy 62.14/54.22 69.17/61.26 75.29/66.31 83.41/74.18 73.21/64.33 86.39/79.41 Semantic Ent. 64.18/56.04 72.29/63.14 77.41/68.22 85.48/76.39 75.17/66.41 88.46/81.27 Lexical Sim. 65.24/57.19 73.33/64.21 78.46/69.37 85.52/77.44 76.31/67.29 88.59/82.31 SelfCheckGPT 58.11/50.28 63.22/55.31 67.38/58.24 74.41/66.33 64.27/56.21 78.46/70.39 RACE 69.14/63.17 76.28/69.41 82.33/74.29 88.47/80.36 80.36/73.22 91.44/85.33 P(true) 67.22/59.26 74.31/66.18 80.41/71.33 87.44/79.28 78.29/69.33 90.38/83.41 FActScore 68.19/61.33 75.39/68.22 81.47/73.38 88.52/81.41 79.34/71.48 91.46/85.37
Long-Context Generation (GovReport).
Table 14: Comparison of methods across different backbone models on GovReport(AUROC/AUPRC).
Method GPT-2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-7B LLaMA2-70B Perplexity 58.13/49.22 64.41/55.37 67.29/58.46 75.34/66.18 63.28/54.33 78.57/69.41 HalluGuard 72.38/66.41 79.27/72.39 84.46/78.31 90.58/84.42 82.44/76.33 93.62/88.51 Inside 69.17/62.24 76.33/68.41 81.44/73.36 88.42/80.31 79.36/71.29 91.47/85.39 MIND 65.21/56.18 72.41/63.29 77.38/68.33 86.33/77.41 75.27/66.38 88.46/82.24 LN-Entropy 63.12/54.27 70.33/61.22 75.41/65.34 83.39/74.21 72.18/62.33 86.38/79.33 Energy 61.09/52.14 69.18/60.41 74.37/64.28 82.34/73.19 71.26/61.44 85.41/78.36 Semantic Ent. 64.17/55.11 71.22/62.31 76.39/67.42 84.46/75.38 73.24/64.28 87.49/80.36 Lexical Sim. 65.26/56.17 72.38/63.38 76.44/68.41 85.43/76.37 74.22/65.19 87.53/81.44 SelfCheckGPT 55.14/46.29 60.31/51.22 63.44/54.19 70.26/60.41 59.33/49.24 73.41/63.38 RACE 68.28/60.33 75.41/66.29 80.36/72.41 87.42/79.33 78.32/70.24 90.38/84.41 P(true) 66.34/57.22 73.39/64.31 78.48/69.44 86.38/77.41 76.33/67.28 89.44/83.36 FActScore 67.41/59.36 74.42/66.41 79.39/71.46 87.41/78.47 77.47/69.44 90.41/84.38
Long-Context Generation (NarrativeQA).
Table 15: Comparison of methods across different backbone models on NarrativeQA(AUROC/AUPRC).
Method GPT-2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-7B LLaMA2-70B Perplexity 56.14/47.22 62.33/53.18 65.41/55.39 72.26/63.41 61.27/51.33 76.38/67.29 HalluGuard 70.36/64.41 77.22/70.37 83.48/76.29 89.53/83.47 81.33/74.36 92.57/87.41 Inside 67.18/60.27 74.39/66.41 80.46/72.31 87.44/79.36 78.41/69.38 90.43/84.32 MIND 63.27/54.18 70.31/61.29 76.33/67.24 84.39/75.41 74.36/64.47 87.41/80.32 LN-Entropy 61.19/52.11 68.27/59.33 73.42/63.21 82.41/73.29 72.14/61.41 85.36/78.44 Energy 60.08/51.14 67.18/58.34 72.37/62.47 81.33/72.41 70.27/60.33 84.44/77.46 Semantic Ent. 63.22/55.09 69.31/61.46 75.44/66.33 83.47/74.41 73.26/63.44 86.47/79.39 Lexical Sim. 64.17/56.22 70.37/62.34 76.41/67.41 84.33/75.44 74.41/65.27 87.46/80.41 SelfCheckGPT 52.14/43.29 57.33/48.31 61.48/51.36 68.41/58.47 56.39/46.31 71.36/61.44 RACE 66.29/58.31 73.42/65.38 79.33/71.28 86.41/78.44 77.28/68.39 89.43/83.38 P(true) 64.31/56.24 71.39/63.33 77.47/68.36 85.38/77.41 75.29/66.33 88.38/82.44 FActScore 65.44/57.36 72.41/64.41 78.52/70.38 86.44/78.33 76.41/68.44 89.44/83.39
@@ Appendix D Usage of LLM
Large language models (LLMs) were employed in a limited and transparent manner during the preparation of this manuscript. Specifically, LLMs were used to assist with linguistic refinement, style adjustments, and minor text editing to improve clarity and readability. They were not involved in formulating the research questions, designing the theoretical framework, conducting experiments, or interpreting results. All scientific contributions-including conceptual development, methodology, analyses, and conclusions-are the sole responsibility of the authors.